A survey of 650 registered voters showed that 360 planned to vote for the candidate of the party that requested the survey. Test the claim that at a(alpha)=0.05 the party's candidate will receive more that 51% of the vote in the election, Use the P-value method.
Answer by stanbon(75887)

A survey of 650 registered voters showed that 360 planned to vote for the candidate of the party that requested the survey.
-----------------------------
Test the claim that at a(alpha)=0.05 the party's candidate will receive more that 51% of the vote in the election, Use the P-value method.
----------------------------------
Ho: p = 0.51
Ha: p > 0.51
-----
sample proportion: 360/650 = 0.55
---
test stat: z(0.55)= (0.55-0.51)/sqrt[0.51*0.49/650] = 2.0400
----
P-value = P(z > 2.400) = 0.0082
=====
Conclusion: Since the p-value is less than 1%, reject Ho at the
1% significance level.
==================================
Cheers,
Stan H.